enrica took two days to drive 1,020 miles. one the first day, she drove 10 hours at an average speed of 62 miles per hour. on the second day, her average speed was 57 miles per hour. how long did she drive on the second day
Answer by richwmiller(17219)

10*62=620 miles first day
1020-620=400 left for second day
400/57=7.01 hours=7hrs 1 min 3 seconds